OpenStudy (jonnyvonny):

Anyone mind illustrating how 1 / (1/2) yields 2? C'mon guys, I don't think its that hard.

OpenStudy (dan815):

1/(1/2) = 1/0.5=2 or 1/1/2 = 1*2/1 = 2

OpenStudy (jonnyvonny):

By illustrate, I mean make pictures, not prove algebraically.

OpenStudy (jonnyvonny):

Should have elaborated, I apologize.

OpenStudy (dan815):

ok

OpenStudy (dan815):

look at a pizza, how many halfs make a full pizza

OpenStudy (dan815):

take a stick, break it in half how many halfs do u have?

OpenStudy (ikram002p):

hehe think of it like this u wanna divide a circle to (halfs (1/2)'s ) ho many half u would have ?